The length of the curve { et cos(t) for 0 <t<l is: y = et sin(t)
Answer #1

Solh find length of come etcosit, y = eutsint OLASI length l= (del, del = √(x²+(41) 2 dit ge = -etsint tetcost y = entlostdel = 2 eht det del = 2. ect det I = va et det [et] da se [ I = 5e [e- eo I = 52 (e-1) A